OpenStudy (anonymous):

Let f(x)= x^2-16 find f^-1(x)

OpenStudy (fibonaccichick666):

so, in general, do you know what the first step to finding an inverse is?

OpenStudy (anonymous):

assuming that x2= x^2 and (f-1)(x) f(x) = x^2 -16 (f-1)(x) = x^2-x-16

OpenStudy (anonymous):

metal?

OpenStudy (fibonaccichick666):

@PlanSlam , please read: http://openstudy.com/code-of-conduct